1. find the gradient to the normal to the curve x^4-4xy+2y^2=0 at point (-1,3)..
2. differentiate f(x)= (x^3-9)^10 with respect of x
1
Expert's answer
2019-09-18T11:53:07-0400

1)We have "F(x,y)=x^4-4xy+2y^2=0".

The implicit function theorem tells us that gradient of the tangent is "y'(x)=-\frac{F'_x(x,y)}{F'_y(x,y)}". The tangent and the normal are perpendicular, so gradient of the normal is "-\frac{1}{y'(x)}=\frac{F'_y(x,y)}{F'_x(x,y)}"

"F'_x(x,y)=4x^3-4y", "F'_y(x,y)=-4x+4y" , so "\frac{1}{y'(x)}=\frac{y-x}{x^3-y}" and "-\frac{1}{y'(-1)}=\frac{3-(-1)}{(-1)^3-3}=-1"

Answer: -1

2)"((x^3-9)^{10})'=10(x^3-9)^9\cdot 3x^2=30x^2(x^3-9)^9"